# Signed Networks: theory, methods, and applications

### v1

@@ 5.1.4 Cascade and threshold models
Threshold models define adoption as a collective process in which a node changes state once the influence of its neighbours exceeds a fixed threshold. In unsigned networks, this influence is additive, so contagion grows monotonically with connectivity. In contrast, in signed networks, negative links reverse or weaken the contribution of active neighbours [srivastavaSocialInfluenceComputation2015, heInformationDiffusionSigned2019]. The impact of these inhibitory effects depends on the network structure [muellerRoleNegativeLinks2023]: scale-free networks stay more resilient because of their hubs, while small-world networks are resilient only to a small fraction of negative links. Negative edges can also represent veto-power interactions [leeThresholdCascadeDynamics2023], where a single antagonistic neighbour prevents activation.
On the other hand, in cascade models, each newly active node has one chance to activate its neighbours with a certain probability, after which it remains active permanently. In signed networks, activation follows the same cascade principle as in unsigned ones, except that negative links invert the influence of active neighbours, causing some nodes to adopt the opposite state of their influencers. A central question is how to choose the initial set of seed nodes to maximise the spread of activation across the network. Determining either the optimal seed set or the final number of activated nodes is a computationally intractable problem in both unsigned and signed networks. A common approach in unsigned networks is to use a greedy algorithm that iteratively selects the most influential node to approximate the optimal seed set. In signed networks, this strategy fails because influence can both increase and decrease through negative links, so the monotonicity required for the greedy approximation guarantee no longer holds. To address this limitation, Srivastava et al. [srivastavaSocialInfluenceComputation2015] propose a heuristic for two competing cascades. Their method selects an inactive node and assigns it a state (+1+1 or 1-1) based on which choice yields the largest expected increase in infections in the next time step. This strategy allows the counterintuitive strategy of seeding some nodes with the opposing opinion to ultimately amplify the target opinion. The authors also extended the algorithm to the linear threshold model.
The Independent Cascade and Linear Threshold models have also been applied to describe information flow between antagonistic communities in social networks. For example, Vendeville et al. [vendevilleModelingEchoChamber2025] used them to show that echo chambers can arise not only in structurally balanced (polarised) networks but also in antibalanced ones. These results challenge the common assumption that structural polarisation and echo‐chamber formation are equivalent.
@@ 5.1.5 Other dynamics
Network synchronisation refers to the process by which the dynamical states of interacting nodes in a network become coordinated over time [arenasSynchronizationComplexNetworks2008] [^26] Some examples of spontaneous synchronisation include neuron activity, chemical oscillators, and fireflies synchronising their flashes.. Each node typically follows its own internal dynamics (e.g., an oscillator, neuron, or agent), but the coupling through the network causes their states to align. In full synchronisation, all nodes evolve identically; in partial synchronisation, only some subsets achieve coordinated behaviour. One of the most paradigmatic models in synchronisation is the Kuramoto model, in which each node $v_{j}$ has an internal phase $\phi_{j}$, which evolves as:
@@
As with other types of dynamics, the presence of negative edges fundamentally alters synchronisation. At the local level, two oscillators connected by a negative link tend to oscillate out of phase. However, if the network is structurally unbalanced, this tendency creates frustration in the collective dynamics. This frustration can lead to slow, glassy relaxation dynamics when the coupling is sufficiently strong. More importantly, the existence of a fully phase-synchronised state (all phases equal) is no longer guaranteed, in contrast with unsigned networks, and one can observe frequency synchronisation without full phase alignment [restrepoSynchronizationLargeDirectedNetworks2006]. Interestingly, under certain conditions, negative links can actually facilitate synchronisation, similar to what happens in consensus dynamics [nishikawaNetworkSynchronizationLandscape2010]. A particularly studied situation involves oscillators known as contrarians, which interact only through negative couplings with the rest of the network [hongKuramotoModelCoupledOscillators2011, hongConformistsContrariansKuramoto2011]. More recent literature focuses on synchronisation in signed bipartite networks, establishing that structural balance is necessary for bipartite synchronisation [zhaiPinningBipartiteSynchronizationAntagonisticSwitching2016, yaghmaieBipartiteCooperativeOutputSynchronizations2017, liuBipartiteSynchronizationLureNetwork2018, liuBipartiteSynchronizationCoupledDelayedNeural2018, xuBipartiteSynchronizationSignedNetworksAperiodicallyIntermittent2021].
Other dynamics explored in signed networks include spiking integrate and fire with excitatory and inhibitory synapses [brunelDynamicsSparselyConnectedExcitatoryInhibitorySpikingNeurons2000, litwinKumarSlowDynamicsHighVariability2012], Hopfield-like dynamics [sompolinskyChaosRandomNeuralNetworks1988, mezardMeanFieldMessagePassingHopfield2016], and Boolean regulatory networks [thomasBooleanFormalizationGeneticControlCircuits1973, remyGraphicRequirementsMultistabilityAttractive2008, richardPositiveNegativeCyclesBoolean2019].
@@ 5.2.1 Toy models of signed networks dynamics
The paradigmatic work in this line of research, i.e., toy models with simple local rules, studied from a theoretical perspective in terms of absorbing states, phase transitions, and long-term behaviour, is the paper by Antal et al. [antalDynamicsSocialBalance2005]. In this model, the authors propose two variants of a triad-based balancing mechanism: Local Triad Dynamics (LTD) and Constrained Triad Dynamics (CTD). At each time step, an unbalanced triangle may flip the sign of one of its links to reach a balanced state. In both variants, if the triangle is of the ()(---) type, it updates to a (+)(+--) triangle with probability one. However, if the triangle is of the (++)(++-) type, it updates to the $(+++)$ type with probability $p$ and to (+)(+--) with probability 1p1-p. In LTD, the updates are always accepted, while in CTD, the updates are only accepted if they do not create new unbalanced triangles elsewhere in the network. Under LTD and in the limit of infinite sizes, the system exhibits a dynamic phase transition: as $p$ varies, the network shifts from a dynamic steady state to a frozen one where all links become positive. In finite systems, the network always reaches a balanced state, whose nature depends on $p$. A similar behaviour is found with CTD, although the dynamics may get trapped in jammed states: configurations that cannot evolve anymore under the update rules, but are not structurally balanced. While earlier models also explored balance dynamics (e.g., [wangSentimentSocialMitosis2003]), Antal et al.’s work remains the most influential example, introducing key ideas for subsequent research like jammed states [marvelEnergyLandscapeSocial2009, chatterjeeSocialBalanceNetworks2022] and balance-unbalance transitions  [antalSocialBalanceNetworks2006, arabzadehLifetimeLinksInfluences2021, rabbaniMeanfieldSolutionStructural2019]. Subsequent extensions have incorporated continuous edge weights [dengStudySignAdjustment2012] and alternative mechanisms, including homophily [gorskiHomophilyBasedFew2020] and status dynamics [gorskiCoexistenceBalanceHierarchies2025]. Finally, some models explore a more sophisticated implementation of a balancing mechanism. For instance, the model proposed in [hummonDynamicsSocialBalance2003] couples Heider’s notion of individual-level balance with Harary’s group-level balance, and explores the interaction between both. Under this model, the system can become frozen in states where local configurations are balanced but the global structure remains unbalanced.
@@
Parallel to the development of Antal et al.’s discrete model, continuous-time approaches also emerged. These models describe the evolution of link weights through differential equations rather than local discrete rules, encoding the drive toward balance in the equation’s structure. A foundational example is [kulakowskiHEIDERBALANCECONTINUOUS2005] (see also [gawronskiHeiderBalanceSocial2005]), which proposes $\frac{dA}{dt}=A^{2}$ as the evolution equation for a fully connected signed network, where $A$ is the adjacency matrix. Simulations show that the system always converges to a balanced state. A more detailed exploration of this model appears in the work of Marvel et al. [marvelContinuoustimeModelStructural2011], who show that, for generic initial conditions, the system always converges either to a fully positive network or to a polarised state with two antagonistic factions. Building on this framework, [summersActiveInfluenceDynamical2013] introduces the idea of targeted influence: the authors show that a single agent can steer the system toward any desired outcome by locally adjusting its friendliness levels, and they provide a method to calculate how much influence is needed to do so.
@@
Given the nature of relationships in a signed network, in which interactions can be hostile or friendly, it is natural to introduce concepts from game theory, such as cooperation and defection, into dynamic models of signed network evolution. The ways to do this are multiple: one can either analyse whether certain games favour the formation of certain structures, or whether certain structures favour the adaptation of players with different strategies.  [traagDynamicalModelsExplaining2013] explored the interplay between balance dynamics and game theory by taking a continuous-time model in which the network evolves driven by different balancing mechanisms. They studied which type of network structures favour a better payoff when the agents play a Prisoner’s Dilemma, assuming that a positive link from $i$ to $j$ represents that $i$ considers $j$ to have a good reputation, and therefore would choose to cooperate with $j$, and a negative link would lead the agent to defect. Another example is [righiEMOTIONALSTRATEGIESCATALYSTS2014], in which agents play a Prisoner’s Dilemma in a signed network, and the network co-evolves with the agents’ strategies (see Section 5.3 for other coevolving models). The authors show that when the strategy depends on the emotional content embedded in the signed network, favourable conditions are created for the emergence of unconditional cooperation (see also [tanEvolutionaryGameApproach2016] and [hillerFriendsEnemiesModel2017]). Finally, [heEvolutionCooperationSigned2018] develops a model in which the coupled dynamics between cooperation and balance are studied. The main difference in this model is that the authors explore the possibility that the signs of the relationships influence the payoffs received in the games considered and study the coupled evolution of strategies and relationships.
@@
Finally, we turn to models that describe networks using a Hamiltonian function. While these models are not strictly dynamic models of network evolution, they incorporate implicit dynamical concepts, as we will see. These models originate from statistical physics and assume that each possible local network configuration (e.g. dyads or triads) is assigned an energy or cost that reflects how likely that local configuration is. This energy is specified by a Hamiltonian function $H(s)$, where $s$ represents the network’s state (e.g., its pattern of connections) [parkStatisticalMechanicsNetworks2004, squartiniAnalyticalMaximumlikelihoodMethod2011]. The model then assumes that configurations with lower energy are more likely, following a probability distribution of the form $P(s)\propto e^{-\beta H(s)}$. The parameter $\beta$ controls how strongly the system favours low-energy states: high $\beta$ enforces strong preference for optimal configurations, while low $\beta$ allows greater randomness [^27] Although here we focus on Hamiltonian approaches to model the dynamics of signed networks, these methods have also played an important role in the signed community detection (see Section 4.4) and in defining null models for signed networks (see Section 4.1). It is worth mentioning that, beyond the context of signed networks specifically, Hamiltonian approaches are closely connected to Exponential Random Graph Models [lusher2013exponential] (see Section 4.1).
This approach was pioneered by Axelrod [axelrodLandscapeTheoryAggregation1993] and Marvel et al. [marvelEnergyLandscapeSocial2009], but one of the most comprehensive overviews can be attributed to Belaza et al. [belazaStatisticalPhysicsBalance2017]. The Hamiltonian they propose is
@@
Where $A$ represents the adjacency matrix. Each element accounts for the cost associated with different dyadic and triadic configurations, weighted by the parameters $\alpha$, $\gamma$, and $\omega$. These parameters determine the influence of each configuration, with $\alpha$ controlling triadic interactions, $\gamma$ the dyadic interactions, and $\omega$ the edge-level contributions. As we can see, the main advantage of this approach is that it provides flexibility, as each term in $H$ can model a different mechanism controlling network dynamics. Another advantage is that the parameters $\alpha,\gamma,\omega$ can be computed from empirical data and used to evaluate the relative importance of the corresponding mechanism. An example is depicted in Figure 15. Other works using Hamiltonian models include [shojaeiPhaseTransitionNetwork2019][belazaSocialStabilityExtended2019][rabbaniMeanfieldSolutionStructural2019] or [manshourDynamicsSocialBalance2021], each introducing different Hamiltonian functions associated with specific phenomena under investigation.
As we mentioned, an important but subtle detail is that these models do not necessarily describe how networks evolve dynamically. Instead, they describe the macroscopic characteristics of the network when it reaches a dynamical equilibrium, an assumption that has been shown to hold in some real networked systems [gonzalez-casadoEvidenceEquilibriumDynamics2025] but cannot be taken for granted. However, the terms in the Hamiltonian include assumptions about the mechanisms that make the networks evolve in time, favouring some configurations over others. Therefore, it can be applied to real systems to extract insights about the implicit dynamics. In fact, once a Hamiltonian is specified, it is possible to define a dynamics that makes the system evolve toward such an equilibrium [landauGuideMonteCarlo2021]. The usual approach is to let the system evolve by randomly proposing small changes (like flipping a link sign) and deciding whether to accept them based on the change in the system’s energy. For example, the Metropolis rule [metropolisEquationStateCalculations1953] always accepts changes that lower the energy of the system, and accepts energy-increasing transitions with a probability $e^{-\beta\Delta E}$ (where $\Delta E$ is the increase in energy).
@@ 5.2.2 Data-driven approaches
One recurring example in the study of signed network dynamics involves interpersonal relations within relatively small groups in closed environments, such as schools, universities, or monasteries. These settings are especially appealing because they make it relatively easy to track relationships over time and collect longitudinal data. The most influential datasets in this area are the Sampson and Nordlie-Newcomb studies. The Sampson dataset [sampsonNovitiatePeriodChange1968] captures the evolution of relations among a group of monks, while the Nordlie-Newcomb dataset [nordlieLONGITUDINALSTUDYINTERPERSONAL1958, newcombAcquaintanceProcess1961] records the evolving relationships among college students from their first encounter over a period of several weeks. These early works laid the empirical foundation for much of the subsequent research on signed network dynamics.
@@
Analyses of the Nordlie-Newcomb dataset (such as those by [nakaoLongitudinalApproachSubgroup1993][doreianBriefHistoryBalance1997], and [doreianPretransitiveBalanceMechanisms2001]) identify three key mechanisms that shape network evolution: reciprocity (tendency for mutual links), transitivity (the tendency to form closed triads), and social balance. These studies show that reciprocity is present early on, while transitivity and balance become increasingly prominent over time. Later analyses also reveal the role of differential popularity, where certain individuals consistently attract more positive or negative links, possibly due to personality traits or group-level perceptions. Differential popularity also shows how mechanisms can interact in misleading ways: it can create an excess of balanced or unbalanced triads, potentially leading researchers to mistakenly infer the presence of a structural balance mechanism if these effects are not properly controlled for. Similar insights emerge from the analysis of the Sampson dataset in [doreianEvolutionHumanSigned2004], which observes a gradual shift toward more balanced, transitive, and reciprocal structures over time.
@@
Apart from these two famous datasets, most of the work on offline social networks has focused on schools, where the structure of students’ personal relationships is analysed. Early studies considered only networks of directed positive links. [katzConceptConfigurationInterpersonal1959] can be considered as a foundational work in this direction, and they analyse for the first time the temporal evolution of such a social network as a Markov chain. This quickly became the state-of-the-art tool to analyse network dynamics [snijdersStatisticalModelsSocial2011]. This work was followed by others which explored the action of different mechanisms at the level of positive links, like transitivity and reciprocity, and their consequences in the network structure [davisStructurePositiveInterpersonal1967, hollandMethodDetectingStructure1970, hollandTransitivityStructuralModels1971, hallinanCommentHollandLeinhardt1972, hallinanStructuralModelSentiment1974, hallinanFriendshipPatternsOpen1976, hallinanDevelopmentChildrensFriendship1977]. The work by Hallinan et al. [hallinanStructuralModelSentiment1974] stands out because it reintroduces the role of the cognitive perspective in the analysis, as Heider suggested (see Section 3.1). Specifically, it develops the idea that, in a network, it is the nodes that experience tension. This idea, which seems obvious, is overlooked, for example, when the analysis is focused on triads, telling us whether they are balanced or unbalanced, and seeing how unbalanced ones can evolve into more balanced states. Hallinan suggests that to understand this evolution, it is necessary to put oneself in the perspective of the nodes, not the triangles, and see how the triad tension affects them. In fact, they use this idea to explain the abundance of certain types of triads in the data that other authors had not been able to explain.
@@
Most studies focusing on schools collect longitudinal data exclusively on positive links. However, a few works have gathered longitudinal network data in schools as closed environments, incorporating also negative links [huitsingVictimsBulliesTheir2014, stadtfeld2020emergence, escribanoEvolutionSocialRelationships2021, escribano2023stability, gonzalez-casadoEvidenceEquilibriumDynamics2025]. Among them, [gonzalez-casadoEvidenceEquilibriumDynamics2025] is particularly relevant, as it shows that the network of social relationships remains in dynamical equilibrium in the statistical–mechanical sense. In other words, while the microscopic dynamics are highly active, the macroscopic properties of the network remain stable. In the context of signed networks, it implies that, even with balancing mechanisms operating, the overall level of balance in the network remains constant despite strong volatility in negative links. Hence, balancing mechanisms coexist and compete with other processes that act to stabilise the level of balance in the system.
@@
Another domain in which signed networks evolution has received attention is international relations. These data are relatively easy to obtain from historical event databases like Correlates of War [palmerMID5Dataset201120142022, giblerInternationalMilitaryAlliances2008] (see Section 6.2). From them, one can reconstruct temporal signed networks where nodes represent states, positive links indicate alliances, and negative links capture tensions, rivalries, or wars.
@@
The first to apply this type of analysis was Harary, who used balance theory to offer a new perspective on conflicts in the Middle East [hararyStructuralAnalysisSituation1961]. A similar analysis was later conducted by Moore [mooreInternationalApplicationHeiders1978, mooreStructuralBalanceInternational1979]. While pioneering, these works have been criticised (e.g. [doreianStructuralBalanceSigned2015]) for relying on arbitrary country selections, short time windows, and limited attention to broader context. They also argued that focusing on particularly acute events may create an illusion of balance that does not reflect broader or long-term dynamics. As an alternative, [doreianStructuralBalanceSigned2015] analysed the international system from 1946 to 1999 using the Correlates of War dataset and found that balance levels fluctuate rather than steadily increase. Later studies, such as [diaz-diazMathematicalModelingLocal2024], examined these fluctuations at the level of individual countries, finding correlations with major historical events occurring in the corresponding countries and years. However, as previously noted, a stable level of balance does not necessarily imply the absence of structural balance as an active mechanism in the system. Its presence is better reflected in the persistent over-representation of certain structures, compared to what would be expected by chance. In this regard, Kirkley et al. [kirkleyBalanceSignedNetworks2019] provided compelling evidence that international relations networks exhibit significant levels of balance.
@@
It should be noted that the analysis of signed networks of international relations is much less developed than in the context of offline social networks. Most existing studies adopt an interpretative perspective, with relatively few efforts aimed at modelling or predicting network evolution. A notable exception is the work by Askarisichani et al. [askarisichani19952018GlobalEvolution2020], who propose a Markov model of triad transitions with demonstrable predictive power. Mechanisms beyond structural balance have also received little attention. In this regard, the international relations literature offers a valuable example of how to critically assess structural balance in relation to alternative theories [healyBalancePowerInternational1973a, maozWhatEnemyMy2007]. There is also a growing need to bridge network-based approaches with traditional geopolitical analysis, for example by combining structural balance theory with factors like geographical proximity or economic ties, and some challenges are still open, such as the inclusion of memory effects in the system (what effect past events of the network have on the present and to what extent they can be modelled as Markovian systems), if the length of cycles plays a role on the balance mechanism, whether link directionality plays a significant role.

### v2

@@ 5.1.4 Cascade and threshold models
Threshold models define adoption as a collective process in which a node changes state once the influence of its neighbours exceeds a fixed threshold. In unsigned networks, this influence is additive, so contagion grows monotonically with connectivity. In contrast, in signed networks, negative links reverse or weaken the contribution of active neighbours [365, 174]. The impact of these inhibitory effects depends on the network structure [281]: scale-free networks stay more resilient because of their hubs, while small-world networks are resilient only to a small fraction of negative links. Negative edges can also represent veto-power interactions [227], where a single antagonistic neighbour prevents activation.
On the other hand, in cascade models, each newly active node has one chance to activate its neighbours with a certain probability, after which it remains active permanently. In signed networks, activation follows the same cascade principle as in unsigned ones, except that negative links invert the influence of active neighbours, causing some nodes to adopt the opposite state of their influencers. A central question is how to choose the initial set of seed nodes to maximise the spread of activation across the network. Determining either the optimal seed set or the final number of activated nodes is a computationally intractable problem in both unsigned and signed networks. A common approach in unsigned networks is to use a greedy algorithm that iteratively selects the most influential node to approximate the optimal seed set. In signed networks, this strategy fails because influence can both increase and decrease through negative links, so the monotonicity required for the greedy approximation guarantee no longer holds. To address this limitation, Srivastava et al. [365] propose a heuristic for two competing cascades. Their method selects an inactive node and assigns it a state (+1+1 or 1-1) based on which choice yields the largest expected increase in infections in the next time step. This strategy allows the counterintuitive strategy of seeding some nodes with the opposing opinion to ultimately amplify the target opinion. The authors also extended the algorithm to the linear threshold model.
The Independent Cascade and Linear Threshold models have also been applied to describe information flow between antagonistic communities in social networks. For example, Vendeville et al. [391] used them to show that echo chambers can arise not only in structurally balanced (polarised) networks but also in antibalanced ones. These results challenge the common assumption that structural polarisation and echo‐chamber formation are equivalent.
@@ 5.1.5 Other dynamics
Network synchronisation refers to the process by which the dynamical states of interacting nodes in a network become coordinated over time [23] [^26] Some examples of spontaneous synchronisation include neuron activity, chemical oscillators, and fireflies synchronising their flashes.. Each node typically follows its own internal dynamics (e.g., an oscillator, neuron, or agent), but the coupling through the network causes their states to align. In full synchronisation, all nodes evolve identically; in partial synchronisation, only some subsets achieve coordinated behaviour. One of the most paradigmatic models in synchronisation is the Kuramoto model, in which each node $v_{j}$ has an internal phase $\phi_{j}$, which evolves as:
@@
As with other types of dynamics, the presence of negative edges fundamentally alters synchronisation. At the local level, two oscillators connected by a negative link tend to oscillate out of phase. However, if the network is structurally unbalanced, this tendency creates frustration in the collective dynamics. This frustration can lead to slow, glassy relaxation dynamics when the coupling is sufficiently strong. More importantly, the existence of a fully phase-synchronised state (all phases equal) is no longer guaranteed, in contrast with unsigned networks, and one can observe frequency synchronisation without full phase alignment [327]. Interestingly, under certain conditions, negative links can actually facilitate synchronisation, similar to what happens in consensus dynamics [297]. A particularly studied situation involves oscillators known as contrarians, which interact only through negative couplings with the rest of the network [184, 183]. More recent literature focuses on synchronisation in signed bipartite networks, establishing that structural balance is necessary for bipartite synchronisation [414, 400, 244, 243, 399].
Other dynamics explored in signed networks include spiking integrate and fire with excitatory and inhibitory synapses [51, 242], Hopfield-like dynamics [359, 270], and Boolean regulatory networks [382, 326, 329].
@@ 5.2.1 Toy models of signed networks dynamics
The paradigmatic work in this line of research, i.e., toy models with simple local rules, studied from a theoretical perspective in terms of absorbing states, phase transitions, and long-term behaviour, is the paper by Antal et al. [15]. In this model, the authors propose two variants of a triad-based balancing mechanism: Local Triad Dynamics (LTD) and Constrained Triad Dynamics (CTD). At each time step, an unbalanced triangle may flip the sign of one of its links to reach a balanced state. In both variants, if the triangle is of the ()(---) type, it updates to a (+)(+--) triangle with probability one. However, if the triangle is of the (++)(++-) type, it updates to the $(+++)$ type with probability $p$ and to (+)(+--) with probability 1p1-p. In LTD, the updates are always accepted, while in CTD, the updates are only accepted if they do not create new unbalanced triangles elsewhere in the network. Under LTD and in the limit of infinite sizes, the system exhibits a dynamic phase transition: as $p$ varies, the network shifts from a dynamic steady state to a frozen one where all links become positive. In finite systems, the network always reaches a balanced state, whose nature depends on $p$. A similar behaviour is found with CTD, although the dynamics may get trapped in jammed states: configurations that cannot evolve anymore under the update rules, but are not structurally balanced. While earlier models also explored balance dynamics (e.g., [394]), Antal et al.’s work remains the most influential example, introducing key ideas for subsequent research like jammed states [262, 63] and balance-unbalance transitions  [16, 17, 322]. Subsequent extensions have incorporated continuous edge weights [83] and alternative mechanisms, including homophily [153] and status dynamics [152]. Finally, some models explore a more sophisticated implementation of a balancing mechanism. For instance, the model proposed in [195] couples Heider’s notion of individual-level balance with Harary’s group-level balance, and explores the interaction between both. Under this model, the system can become frozen in states where local configurations are balanced, but the global structure remains unbalanced.
@@
Parallel to the development of Antal et al.’s discrete model, continuous-time approaches also emerged. These models describe the evolution of link weights through differential equations rather than local discrete rules, encoding the drive toward balance in the equation’s structure. A foundational example is [220] (see also [144]), which proposes $\frac{dA}{dt}=A^{2}$ as the evolution equation for a fully connected signed network, where $A$ is the adjacency matrix. Simulations show that the system always converges to a balanced state. A more detailed exploration of this model appears in the work of Marvel et al. [261], who show that, for generic initial conditions, the system always converges either to a fully positive network or to a polarised state with two antagonistic factions. Building on this framework, [369] introduces the idea of targeted influence: the authors show that a single agent can steer the system toward any desired outcome by locally adjusting its friendliness levels, and they provide a method to calculate how much influence is needed to do so.
@@
Given the nature of relationships in a signed network, in which interactions can be hostile or friendly, it is natural to introduce concepts from game theory, such as cooperation and defection, into dynamic models of signed network evolution. The ways to do this are multiple: one can either analyse whether certain games favour the formation of certain structures, or whether certain structures favour the adaptation of players with different strategies.  [389] explored the interplay between balance dynamics and game theory by taking a continuous-time model in which the network evolves driven by different balancing mechanisms. They studied which type of network structures favour a better payoff when the agents play a Prisoner’s Dilemma, assuming that a positive link from $i$ to $j$ represents that $i$ considers $j$ to have a good reputation, and therefore would choose to cooperate with $j$, and a negative link would lead the agent to defect. Another example is [330], in which agents play a Prisoner’s Dilemma in a signed network, and the network co-evolves with the agents’ strategies (see Section 5.3 for other coevolving models). The authors show that when the strategy depends on the emotional content embedded in the signed network, favourable conditions are created for the emergence of unconditional cooperation (see also [377] and [178]). Finally, [173] develops a model in which the coupled dynamics between cooperation and balance are studied. The main difference in this model is that the authors explore the possibility that the signs of the relationships influence the payoffs received in the games considered and study the coupled evolution of strategies and relationships.
@@
Finally, we turn to models that describe networks using a Hamiltonian function. While these models are not strictly dynamic models of network evolution, they incorporate implicit dynamical concepts, as we will see. These models originate from statistical physics and assume that each possible local network configuration (e.g., dyads or triads) is assigned an energy or cost that reflects how likely that local configuration is. This energy is specified by a Hamiltonian function $H(s)$, where $s$ represents the network’s state (e.g., its pattern of connections) [304, 364]. The model then assumes that configurations with lower energy are more likely, following a probability distribution of the form $P(s)\propto e^{-\beta H(s)}$. The parameter $\beta$ controls how strongly the system favours low-energy states: high $\beta$ enforces strong preference for optimal configurations, while low $\beta$ allows greater randomness [^27] Although here we focus on Hamiltonian approaches to model the dynamics of signed networks, these methods have also played an important role in the signed community detection (see Section 4.4) and in defining null models for signed networks (see Section 4.1). It is worth mentioning that, beyond the context of signed networks specifically, Hamiltonian approaches are closely connected to Exponential Random Graph Models [252] (see Section 4.1).
This approach was pioneered by Axelrod [27] and Marvel et al. [262], but one of the most comprehensive overviews can be attributed to Belaza et al. [35]. The Hamiltonian they propose is
@@
Where $A$ represents the adjacency matrix. Each element accounts for the cost associated with different dyadic and triadic configurations, weighted by the parameters $\alpha$, $\gamma$, and $\omega$. These parameters determine the influence of each configuration, with $\alpha$ controlling triadic interactions, $\gamma$ the dyadic interactions, and $\omega$ the edge-level contributions. As we can see, the main advantage of this approach is that it provides flexibility, as each term in $H$ can model a different mechanism controlling network dynamics. Another advantage is that the parameters $\alpha,\gamma,\omega$ can be computed from empirical data and used to evaluate the relative importance of the corresponding mechanism. An example is depicted in Figure 15. Other works using Hamiltonian models include [355][36][322] or [257], each introducing different Hamiltonian functions associated with specific phenomena under investigation.
As we mentioned, an important but subtle detail is that these models do not necessarily describe how networks evolve dynamically. Instead, they describe the macroscopic characteristics of the network when it reaches a dynamical equilibrium, an assumption that has been shown to hold in some real networked systems [151] but cannot be taken for granted. However, the terms in the Hamiltonian include assumptions about the mechanisms that make the networks evolve in time, favouring some configurations over others. Therefore, it can be applied to real systems to extract insights about the implicit dynamics. In fact, once a Hamiltonian is specified, it is possible to define a dynamics that makes the system evolve toward such an equilibrium [226]. The usual approach is to let the system evolve by randomly proposing small changes (like flipping a link sign) and deciding whether to accept them based on the change in the system’s energy. For example, the Metropolis rule [268] always accepts changes that lower the energy of the system, and accepts energy-increasing transitions with a probability $e^{-\beta\Delta E}$ (where $\Delta E$ is the increase in energy).
@@ 5.2.2 Data-driven approaches
One recurring example in the study of signed network dynamics involves interpersonal relations within relatively small groups in closed environments, such as schools, universities, or monasteries. These settings are especially appealing because they make it relatively easy to track relationships over time and collect longitudinal data. The most influential datasets in this area are the Sampson and Nordlie-Newcomb studies. The Sampson dataset [339] captures the evolution of relations among a group of monks, while the Nordlie-Newcomb dataset [298, 292] records the evolving relationships among college students from their first encounter over a period of several weeks. These early works laid the empirical foundation for much of the subsequent research on signed network dynamics.
@@
Analyses of the Nordlie-Newcomb dataset (such as those by [283][94], and [95]) identify three key mechanisms that shape network evolution: reciprocity (tendency for mutual links), transitivity (the tendency to form closed triads), and social balance. These studies show that reciprocity is present early on, while transitivity and balance become increasingly prominent over time. Later analyses also reveal the role of differential popularity, where certain individuals consistently attract more positive or negative links, possibly due to personality traits or group-level perceptions. Differential popularity also shows how mechanisms can interact in misleading ways: it can create an excess of balanced or unbalanced triads, potentially leading researchers to mistakenly infer the presence of a structural balance mechanism if these effects are not properly controlled for. Similar insights emerge from the analysis of the Sampson dataset in [100], which observes a gradual shift toward more balanced, transitive, and reciprocal structures over time.
@@
Apart from these two famous datasets, most of the work on offline social networks has focused on schools, where the structure of students’ personal relationships is analysed. Early studies considered only networks of directed positive links. [206] can be considered as a foundational work in this direction, and they analyse for the first time the temporal evolution of such a social network as a Markov chain. This quickly became the state-of-the-art tool to analyse network dynamics [358]. This work was followed by others which explored the action of different mechanisms at the level of positive links, like transitivity and reciprocity, and their consequences in the network structure [78, 181, 182, 165, 162, 163, 164]. The work by Hallinan et al. [162] stands out because it reintroduces the role of the cognitive perspective in the analysis, as Heider suggested (see Section 3.1). Specifically, it develops the idea that, in a network, it is the nodes that experience tension. This idea, which seems obvious, is overlooked, for example, when the analysis is focused on triads, telling us whether they are balanced or unbalanced, and seeing how unbalanced ones can evolve into more balanced states. Hallinan suggests that to understand this evolution, it is necessary to put oneself in the perspective of the nodes, not the triangles, and see how the triad tension affects them. In fact, they use this idea to explain the abundance of certain types of triads in the data that other authors had not been able to explain.
@@
Most studies focusing on schools collect longitudinal data exclusively on positive links. However, a few works have gathered longitudinal network data in schools as closed environments, incorporating also negative links [194, 366, 103, 104, 151]. Among them, [151] is particularly relevant, as it shows that the network of social relationships remains in dynamical equilibrium in the statistical–mechanical sense. In other words, while the microscopic dynamics are highly active, the macroscopic properties of the network remain stable. In the context of signed networks, it implies that, even with balancing mechanisms operating, the overall level of balance in the network remains constant despite strong volatility in negative links. Hence, balancing mechanisms coexist and compete with other processes that act to stabilise the level of balance in the system.
@@
Another domain in which signed networks evolution has received attention is international relations. These data are relatively easy to obtain from historical event databases like Correlates of War [301, 145] (see Section 6.2). From them, one can reconstruct temporal signed networks where nodes represent states, positive links indicate alliances, and negative links capture tensions, rivalries, or wars.
@@
The first to apply this type of analysis was Harary, who used balance theory to offer a new perspective on conflicts in the Middle East [172]. A similar analysis was later conducted by Moore [277, 278]. While pioneering, these works have been criticised (e.g. [99]) for relying on arbitrary country selections, short time windows, and limited attention to broader context. They also argued that focusing on particularly acute events may create an illusion of balance that does not reflect broader or long-term dynamics. As an alternative, [99] analysed the international system from 1946 to 1999 using the Correlates of War dataset and found that balance levels fluctuate rather than steadily increase. Later studies, such as [90], examined these fluctuations at the level of individual countries, finding correlations with major historical events occurring in the corresponding countries and years. However, as previously noted, a stable level of balance does not necessarily imply the absence of structural balance as an active mechanism in the system. Its presence is better reflected in the persistent over-representation of certain structures, compared to what would be expected by chance. In this regard, Kirkley et al. [212] provided compelling evidence that international relations networks exhibit significant levels of balance.
@@
It should be noted that the analysis of signed networks of international relations is much less developed than in the context of offline social networks. Most existing studies adopt an interpretative perspective, with relatively few efforts aimed at modelling or predicting network evolution. A notable exception is the work by Askarisichani et al. [26], who propose a Markov model of triad transitions with demonstrable predictive power. Mechanisms beyond structural balance have also received little attention. In this regard, the international relations literature offers a valuable example of how to critically assess structural balance in relation to alternative theories [176, 258]. There is also a growing need to bridge network-based approaches with traditional geopolitical analysis, for example by combining structural balance theory with factors like geographical proximity or economic ties, and some challenges are still open, such as the inclusion of memory effects in the system (what effect past events of the network have on the present and to what extent they can be modelled as Markovian systems), if the length of cycles plays a role on the balance mechanism, whether link directionality plays a significant role.